# Merging LoRAs like Playing LEGO: Pushing the Modularity of LoRA to Extremes Through Rank-Wise Clustering

### v1

@@ Appendix B Connection with Vanilla LoRA Composition Methods
The vanilla LoRA composition comprises two training-free methods: the Ensemble of LoRAs and the Fusion of LoRAs. The Ensemble of LoRAs strategy involves the aggregation of the outputs of each submodule within the assembled LoRAs. Let us denote ${\mathbb{A}}=\{{\bm{A}}_{1},{\bm{A}}_{2},\ldots,{\bm{A}}_{n}\}$ and ${\mathbb{B}}=\{{\bm{B}}_{1},{\bm{B}}_{2},\ldots,{\bm{B}}_{n}\}$ as the sets representing submodules within $n$ LoRAs. For an input ${\bm{x}}_{i}$, the output derived from the Ensemble of LoRAs can be expressed as ${\bm{x}}_{i}^{\prime}=\frac{1}{n}\sum_{j=1}^{n}{\bm{B}}_{j}{\bm{A}}_{j}{\bm{x}}_{i}$, where ${\bm{x}}_{i}^{\prime}$ denotes the output.
In contrast to the Ensemble method, which combines the output of different LoRAs, model merging presents an alternative composition strategy. Let the parameters of each LoRA, $\phi_{i}$, be denoted by $\Theta_{i}$. A typical strategy involves averaging these parameters, represented as $\Theta_{\text{fusion}}=\frac{1}{k}\sum_{j=1}^{k}\Theta_{j}$. This formulation allows the fused parameter to function akin to a single LoRA.
Our proposed LoRA-LEGO method offers refined control over model complexity and performance, facilitating an optimal balance between semantic integration and computational efficiency. Unlike Simple Fusion, which lacks alignment and may introduce inconsistencies across merged LoRAs, our approach aligns and selectively fuses MSUs based on their semantic closeness. This selective fusion significantly reduces redundancy, a prevalent issue in the Ensemble method where all MSUs are treated uniformly, potentially leading to an excessive parameter count that can degrade performance. Our clustering-based strategy addresses these challenges by effectively condensing the most relevant semantic features into fewer clusters, thereby enhancing both the model’s performance and its inference efficiency.
@@ Appendix C Details of Baseline Methods
    Weight Averaging. This approach averages the parameters across different instances of LoRA, resulting in a new composite LoRA defined as ${\bm{A}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{A}}_{i}$ and ${\bm{B}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{B}}_{i}$, where ${\bm{A}}_{i}$ and $B_{i}$ represent the parameters from the $i$-th instance of the original LoRA models, and $n$ is the number of models being averaged.

### v3

@@ Appendix B Connection with Vanilla LoRA Composition Methods
The vanilla LoRA composition can be categoried into two types of training-free methods: the model ensembling and model merging Tang et al. (2024). The ensemble strategy involves aggregating the outputs of each submodule within the assembled LoRAs. Let us denote $\mathcal{A}=\{{\bm{A}}_{1},{\bm{A}}_{2},\ldots,{\bm{A}}_{n}\}$ and =$\mathcal{B}=\{{\bm{B}}_{1},{\bm{B}}_{2},\ldots,{\bm{B}}_{n}\}$ as the sets representing submodules within $n$ LoRAs. For an input ${\bm{x}}_{i}$, the output derived from the ensemble of LoRAs can be expressed as ${\bm{x}}_{i}^{\prime}=\frac{1}{n}\sum_{j=1}^{n}{\bm{B}}_{j}{\bm{A}}_{j}{\bm{x}}_{i}$, where ${\bm{x}}_{i}^{\prime}$ denotes the output. The performance of the ensemble of LoRAs tends to be more stable, but it incurs additional computational overhead. In contrast to the ensemble method, model merging presents an alternative composition strategy. A typical strategy involves employing element-wise fusion of these parameters, represented as ${\bm{A}}^{\prime}=\frac{1}{n}\sum_{j=1}^{n}{\bm{A}}_{j}$ and ${\bm{B}}^{\prime}=\frac{1}{n}\sum_{j=1}^{n}{\bm{B}}_{j}$. This formulation allows the merged parameters to function similarly to a single LoRA. However, directly merging parameters can lead to performance degradation due to parameter interference.
Our proposed LoRA-LEGO method serves as a bridge between the two strategies, ensuring an optimal balance between computational efficiency and performance. By selectively aligning and fusing MSUs based on their semantic similarity, LoRA-LEGO effectively condenses the most relevant semantic features into fewer clusters. This process allows for the merging of parameters within each cluster, reducing the overall parameter count in a manner similar to the model merging method. By adjusting the number of clusters, LoRA-LEGO can accommodate more parameters for inference, much like the ensemble method. In this way, our method leverages the strengths of both methodologies, ultimately enhancing model performance and inference efficiency.
@@ Appendix C Details of Baseline Methods
    Weight Averaging. This approach averages the parameters across different instances of LoRA, resulting in a new composite LoRA defined as ${\bm{A}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{A}}_{i}$ and ${\bm{B}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{B}}_{i}$, where ${\bm{A}}_{i}$ and ${\bm{B}}_{i}$ represent the parameters from the $i$-th instance of the original LoRA models, and $n$ is the number of models being averaged.